OpenStudy (anonymous):

Prove for every integer (n), n^5-n = 0 (mod 30).

OpenStudy (tkhunny):

Have you considered factoring?

OpenStudy (tkhunny):

Factor it.

OpenStudy (tkhunny):

Why? Factor it.

OpenStudy (anonymous):

@zepdrix

OpenStudy (ikram002p):

ohk ?

OpenStudy (anonymous):

hihihi, I tried induction, but get stuck

OpenStudy (ikram002p):

n^5-n = 0 (mod 30). u can use induction also another way if u wanna use mod n(n^4-1)=0 (mod 30) chose one to start :D

OpenStudy (anonymous):

for hypothesis step, I got (k-1)k(k+1)(k^2+1) but the last step is k(k+1)(k+2)((k+1)^2+1) then stuck, hihi

OpenStudy (anonymous):

either way. Greedy me, I want to know both. hihi

OpenStudy (ikram002p):

mmm well lets induction it for n=1 it work lets assume for n n^5-n = 0 (mod 30). n( n^4-1) = 0 (mod 30). n( n+1)(n-1)(n^2+1) = 0 (mod 30). is true prove for n+1 (n+1)((n+1)^2-1) ((n+1)^2+1) lol im not getting the induction step xD

OpenStudy (anonymous):

neither I, That's why I tagged others to solve it. haiz. it's not a piece of cake.

OpenStudy (ikram002p):

xD

OpenStudy (ikram002p):

lets dump the induction and try some number theory

OpenStudy (ikram002p):

n( n^4-1) = 0 (mod 30). if this one is true then .. n( n^4-1) = 0 (mod 2*3*5).

OpenStudy (anonymous):

It's up to you . I don't know number theory.:)

OpenStudy (ikram002p):

ok did u got that step ?

OpenStudy (anonymous):

among the hypothesis, we have (n-1)(n)(n+1) , they are 3 consecutive integer, so that they divided by 2 for sure :)

OpenStudy (ikram002p):

so we have 3 equtions n( n^4-1) = 0 (mod 2 ). n( n^4-1) = 0 (mod 3). n( n^4-1) = 0 (mod 5).

OpenStudy (ikram002p):

good u already proved it for 2 :D

OpenStudy (ikram002p):

nw for 3 any number would devide 3 should be of the form 3k right ?

OpenStudy (ikram002p):

using devisition algorithm n have 3 choices n=3m n=3m+1 n=3m+2

OpenStudy (ikram002p):

got it tell nw ?

OpenStudy (anonymous):

let finish the proof. I will ask if I have question

OpenStudy (anonymous):

got it

OpenStudy (ikram002p):

so for n=3m n( n^4-1) = 0 (mod 3 ) 3m(....)=0(mod 3 ) devide 3 for n=3m+1 (3m+1)(3m+2)(3m) ((3m)^2+1)) = 0 (mod 3 ) as u can see there is 3m so devide 3

OpenStudy (ikram002p):

for (3m+2) n( n+1)(n-1)(n^2+1) (3m+2)(3m+3) (....)=0mod 30 3(...) =0mod 30 so it devide 3

OpenStudy (anonymous):

I got this part (prove the expression divided by 3) how about 5

OpenStudy (ikram002p):

ok about 5 mmm the numer devide 5 when first digit =0,5 but that wont help in proof

OpenStudy (ikram002p):

other case the number devide 5 when its of the form 5G we should use D.A again :D for g=5 so n would have 5 cases n=5k n=5k+1 n=5k+2 n=5k+3 n=5k+4

OpenStudy (anonymous):

ok, I got 5 too. Thank you hihihi

OpenStudy (anonymous):

However, my question is about how to combine them?

OpenStudy (ikram002p):

lolz whats hihihih :P

OpenStudy (ikram002p):

combine what ?

OpenStudy (anonymous):

since the proof of 2 include in 3, (3 consecutive numbers) but for 5, we assume that n = 5k,..... which is not include 2 and 3 in

OpenStudy (ikram002p):

well its a theorm that set of \( d_0 ,d_1, d_2 ...\) are devisors of a number g and n = 0 mod g then n= 0 mod \( [d_0×d_1× d_2×...]\)

OpenStudy (ikram002p):

i dint got ur question , do u wanna prove it for 5 ?

OpenStudy (anonymous):

Yes, but n is vivid you set n =5m , and we have 5m divided by 5, but how it divided by 3 also?

OpenStudy (ikram002p):

O.O

OpenStudy (ikram002p):

u said u got that part

OpenStudy (anonymous):

I am sorry for my bad English to interpret my mind. let me arrange it for 3, you set n =3m, n = 3m+1, n=3m+2 and our expression: (n-1)(n)(n+1) (n^2+1) , now replace n = 3m for example, we still have 3m (3m -1)(3m+1) (this part is 3 consecutive numbers also, Therefore, this part divided by 2 also.That's what I mean divided by 2 is include the part of divided by 3) but divided by 5 is another problem. It included 2 part, but not 3 .

OpenStudy (ikram002p):

mmm lol ok u dint get the concept of D.A when i set this n^5-n = 0 (mod 3 ). then forget about 2 or 5 im only trying to prove for 3

OpenStudy (ikram002p):

and since u dnt know this concept i suggest not to continue ... and try something else

OpenStudy (ikram002p):

what is your course name ?

OpenStudy (anonymous):

can you give me a link to learn more ?

OpenStudy (ikram002p):

i might find some other method

OpenStudy (ikram002p):

yeah sure

OpenStudy (ikram002p):

wait

OpenStudy (anonymous):

hihihi.... it's not my problem. it's my curiosity.

OpenStudy (ikram002p):

wow ! where did u find it ? how old are you ?

OpenStudy (anonymous):

this post is not mine, either. hihihj

OpenStudy (ikram002p):

ohh ic ! well ull gonna find it easy !

OpenStudy (ikram002p):

here , go through chapter 1 and 2 https://drive.google.com/file/d/0B1oczGnMVhNqTDMtTEs0QzFPTm8/edit?usp=sharing

OpenStudy (ikram002p):

let me know if u got it , or if u need something else :D

OpenStudy (anonymous):

thank you

OpenStudy (ikram002p):

np :)

OpenStudy (tkhunny):

\(n\cdot(n-1)\cdot(n+1)\cdot (n^{2}+1)\) Obviously, with three consecutive integers, n-1, n, and n+1, this thing is divisible by 2 and by 3. The only challenge is 5. Can we prove that either \(n^{3} + n\) or \(n^{3} - n\) is divisible by 5?

OpenStudy (ikram002p):

@tkhunny there are two ways to prove it in number theory 1_ using difision algorithm for 5and try 5 possible values n=5m n=5n+1 n=5n+2 n=5m+3 n=5m+4 n=5m+5 2_there is a theorim Euler that sit for any number a and prime P , gcd(a,p)=1 then a^5=a mod 5 so a^5-a=0 mod 5 done